# BioCLIP: A Vision Foundation Model for the Tree of Life

### v2

@@ Abstract
Images of the natural world, collected by a variety of cameras, from drones to individual phones, are increasingly abundant sources of biological information. There is an explosion of computational methods and tools, particularly computer vision, for extracting biologically relevant information from images for science and conservation. Yet most of these are bespoke approaches designed for a specific task and are not easily adaptable or extendable to new questions, contexts, and datasets. A vision model for general organismal biology questions on images is of timely need. To approach this, we curate and release TreeOfLife-10M, the largest and most diverse ML-ready dataset of biology images. We then develop BioCLIP, a foundation model for the tree of life, leveraging the unique properties of biology captured by TreeOfLife-10M, namely the abundance and variety of images of plants, animals, and fungi, together with the availability of rich structured biological knowledge. We rigorously benchmark our approach on diverse fine-grained biology classification tasks, and find that BioCLIP consistently and substantially outperforms existing baselines (by 17% to 20% absolute). Intrinsic evaluation reveals that BioCLIP has learned a hierarchical representation conforming to the tree of life, shedding light on its strong generalizability. [^1] See github.com/Imageomics/bioclip for models, data and code.
@@ 1 Introduction
However, applying computer vision to answer a biological question is still a laborious task requiring substantial machine learning expertise and effort—biologists must manually label sufficient data for the specific taxa and task of interest, and find and train a suitable model for the task. Meanwhile, foundation models [12] such as CLIP [66] and GPT-3 [14] are extraordinarily valuable by enabling zero-shot or few-shot learning for a wide range of tasks. An analogous vision foundation model for biology should be useful for tasks spanning the entire tree of life [34, 50] instead of just the taxa it has been trained on. That would significantly lower the barrier to apply AI to biology.
@@
While the goals of generalization, fine-grained classification, and data efficiency are not new in computer vision, existing general-domain vision models [66, 92, 58] trained on hundreds of millions of images fall short when applied to evolutionary biology and ecology. Specifically, existing vision models produce general fine-grained representations, useful for comparing common organisms like dogs and wolves, but not for more fine-grained comparisons, e.g., Onoclea sensibilis L. and Onoclea hintonii (see Fig. 1).
@@
1. TreeOfLife-10M: a large-scale, diverse ML-ready biology image dataset. We curate and release the largest-to-date ML-ready dataset of biology images with associated taxonomic labels, containing over $10$ million images covering $454$ thousand taxa in the tree of life. [^2] By ML-ready, we mean the data is standardized in a format suitable for training ML models and is readily available for downloading. In comparison, the current largest ML-ready biology image dataset, iNat21 [82], contains only $2.7$ million images covering $10$ thousand taxa. TreeOfLife-10M integrates existing high-quality datasets like iNat21 and Bioscan-1M [26]. More importantly, it includes newly curated images from the Encyclopedia of Life (eol.org), which supplies most of TreeOfLife-10M’s data diversity. Every image in TreeOfLife-10M is labeled with its taxonomic hierarchy to the finest level possible, as well as higher taxonomic ranks in the tree of life (see Figs. 1 and 3 for examples of taxonomic ranks and labels). TreeOfLife-10M enables training BioCLIP and future biology foundation models.
2. BioCLIP: a vision foundation model for the tree of life. With a large-scale labeled dataset like TreeOfLife-10M, a standard, intuitive training strategy (as adopted by other vision models like ResNet50 [31] and Swin Transformer [45]) is to use a supervised classification objective and learn to predict the taxonomic indices from an image. However, this fails to recognize and leverage the rich structure of taxonomic labels—taxa do not exist in isolation but are interconnected in a comprehensive taxonomy. Consequently, a model trained via plain supervised classification may not generalize well to taxa unseen in training, nor could it support zero-shot classification of unseen taxa.
@@
3. Comprehensive benchmarking. We comprehensively evaluate BioCLIP on $10$ fine-grained image classification datasets covering animals, plants, and fungi, including a newly curated Rare Species dataset unseen in training. BioCLIP achieves strong performance in both zero-shot and few-shot settings and substantially outperforms both CLIP [66] and OpenCLIP [39], leading to an average absolute improvement of 18% (zero-shot) and 17% (few-shot). Intrinsic analysis further reveals that BioCLIP has learned a more fine-grained hierarchical representation conforming to the tree of life, explaining its superior generalization.
@@ 2 TreeOfLife-10M
Recent work has shown that data quality and diversity is critical when training CLIP models [22, 54, 24]. We curate TreeOfLife-10M, the most diverse large-scale public ML-ready dataset for computer vision models in biology.
@@ 2.1 Images
The largest ML-ready biology image dataset is iNat21 [82], which contains 2.72.7M images of 1010K species. Despite this class breadth compared to popular general-domain datasets like ImageNet-1K [67], 1010K species are rather limited in the context of biology. The International Union for Conservation of Nature (IUCN) reported over 22M total described species in 2022, with over 1010K bird species and over 1010K reptile species alone [41]. iNat21’s species diversity limits its potential for pre-training a foundation model for the entire tree of life.
@@ 2.2 Metadata & Aggregation
The TreeOfLife-10M dataset integrates iNat21 (training split), our curated EOL dataset, and Bioscan-1M by aggregating the images and canonicalizing the labels. This is a highly non-trivial task because taxonomic hierarchies are notoriously noisy and rarely consistent between sources [33, 29, 49, 60, 4], likely contributing to the prior lack of image datasets large enough to train a foundation-scale vision model for the entire tree of life. We carefully unify and backfill taxonomic hierarchies from EOL, the Integrated Taxonomic Information System (ITIS) [40] and iNaturalist with special consideration for the existence of homonyms (genus-species labels shared among higher-order taxa). For more information on this process, the challenges, our solutions, and remaining issues, see Appendix C.
@@ 2.3 Release & Statistics
We will release our curated training and test data (TreeOfLife-10M and Rare Species, described in Sec. 4.2) on Hugging Face (with DOIs) under a public domain waiver, to the extent primary source licenses allow. This will include a CSV with image metadata and links to the primary sources, accompanied by a GitHub repository with the scripts to generate the datasets. [^4] We encourage future work to cite iNat21 [82], Bioscan-1M [26] and to appropriately attribute images from EOL based on their licenses if citing TreeOfLife-10M.
@@ 3.1 Why CLIP?
CLIP trains two uni-modal embedding models, a vision encoder and a text encoder, to (1) maximize feature similarity between positive (image, text) pairs and (2) minimize feature similarity between negative (image, text) pairs, where positive pairs from the training data and negative pairs are all other possible (image, text) pairings in a batch. After training, CLIP’s encoder models embed individual instances of their respective modalities into a shared feature space. Next, we discuss formatting the text input to CLIP to incorporate the taxonomic structure.
@@ 4.1 Training and Evaluation Details
To train BioCLIP, we initialize from OpenAI’s CLIP weights [66] with a ViT-B/16 vision transformer [20] image encoder and a $77$-token causal autoregressive transformer text encoder. We continue pre-training on TreeOfLife-10M for 100 epochs with a cosine learning rate schedule [46]. We train on $8$ NVIDIA A100-80GB GPUs over $2$ nodes with a global batch size of $32,768$. We also train a baseline model on only the iNat21 dataset and multiple ablation models on 1M examples randomly sampled from TreeOfLife-10M (Secs. 4.3 and 4.4), following the same procedure for BioCLIP except with a smaller global batch size of $16,384$ on $4$ NVIDIA A100 GPUs on $1$ node. All hyperparameters and training details are in Appendix D; the training and evaluation code will be released.
@@
We evaluate on 10 different classification tasks: the 8 biologically-relevant tasks from Meta-Album [80], Birds 525 [65] and our new Rare Species task (described in Sec. 4.2). Meta-Album is a dataset collection for meta-learning, encompassing various subjects. Specifically, we use the Plankton, Insects, Insects 2, PlantNet, Fungi, PlantVillage, Medicinal Leaf, and PlantDoc datasets. Our classification tasks cover all four multi-celled kingdoms in the tree of life (animals, plants, fungi, and protists) and have a diverse image distribution (photographs, microscope images, drawings, and museum specimens). Tab. 2 provides an overview of the datasets; they comprise a variety of label types from full taxonomic names to only scientific or common name, or a mix of the latter.
@@
For zero-shot learning, we follow the same procedure as CLIP. For few-shot learning, we follow SimpleShot [87] to use a nearest-centroid classifier. For $k$-shot learning, we first randomly sample $k$ examples for each class and obtain the image embedding from the visual encoder of the pre-trained models. We then compute the average feature vector of the $k$ embeddings as the centroid for each class. All the examples left in the dataset are used for testing. After applying mean subtraction and L2-normalization to each centroid and test feature vector, we choose the class with the nearest centroid to the test vector as the prediction. We repeat each few-shot experiment $5$ times with different random seeds and report the mean accuracy in Tab. 4. Results with standard deviations are reported in Appendix E.
@@
We compare BioCLIP with the original OpenAI CLIP [66] as well as OpenCLIP [39] trained on LAION-400M [70]. Intuitively, common names of organisms are most pervasive in the training data of CLIP and OpenCLIP and these models work best with common names. This is also confirmed in our preliminary tests. Therefore, we use common names as class labels for CLIP and OpenCLIP by default unless unavailable for a dataset. BioCLIP can leverage taxonomic names, so we use taxonomic + common names by default. However, as noted in Tab. 2, the test datasets come in a variety of labels. Whenever the preferred label type is not available, we use labels that come with the dataset.
@@
CLIP $49.9$ $3.2$ $9.1$ $9.8$ $58.5$ $10.2$ $5.4$ $15.9$ $26.1$ $26.6$ $21.4$
OpenCLIP $54.7$ $2.2$ $6.5$ $9.6$ $50.2$ $5.7$ $8.0$ $12.4$ $25.8$ $31.0$ $20.6$ 0.8-0.8
BioCLIP $72.1$ $6.1$ $34.8$ $20.4$ $91.4$ $40.7$ $24.4$ $38.6$ $28.4$ $37.8$ $39.4$ +18.0+18.0
– iNat21 Only $56.1$ $2.6$ $30.7$ $11.5$ $88.2$ $43.0$ $18.4$ $25.6$ $20.5$ $19.4$ $31.6$ +10.2+10.2
@@
CLIP $43.7$ $25.1$ $21.6$ $13.7$ $42.1$ $17.2$ $49.7$ $70.1$ $24.8$ $28.4$ $33.6$
OpenCLIP $53.7$ $32.3$ $23.2$ $14.3$ $45.1$ $18.4$ $53.6$ $71.2$ $26.8$ $29.3$ $36.7$ +3.1+3.1
BioCLIP $71.8$ $30.6$ $57.4$ $20.4$ $64.5$ $40.3$ $58.8$ $84.3$ $30.7$ $45.3$ $50.4$ +16.8+16.8
– iNat21 Only $74.8$ $29.6$ $53.9$ $19.7$ $67.4$ $35.5$ $55.2$ $75.1$ $27.8$ $37.6$ $47.6$ +14.0+14.0
@@
CLIP $73.5$ $41.2$ $39.9$ $24.6$ $65.2$ $27.9$ $71.8$ $89.7$ $35.2$ $46.9$ $51.5$
OpenCLIP $81.9$ $52.5$ $42.6$ $25.0$ $68.0$ $30.6$ $77.8$ $91.3$ $42.0$ $48.4$ $56.0$ +4.5+4.5
BioCLIP $90.0$ $49.3$ $77.8$ $33.6$ $85.6$ $62.3$ $80.9$ $95.9$ $47.5$ $66.4$ $68.9$ +17.4+17.4
– iNat21 Only $90.1$ $48.2$ $73.7$ $32.1$ $84.7$ $55.6$ $77.2$ $93.5$ $41.0$ $56.0$ $65.1$ +13.6+13.6
Table 4: Zero-, one- and five-shot classification top-1 accuracy for different CLIP models. Bold indicates best accuracy. All models use the same architecture: ViT-B/16 vision encoders, 77-token text encoder. “iNat21 Only” follows the same procedure as BioCLIP but uses iNat21 instead of TreeOfLife-10M. $\Delta$ denotes the difference in mean accuracy with CLIP.
@@
ToL-1M Com $25.5$ $9.5$ $10.6$ $22.6$ $23.1$
Sci $11.1$ $22.2$ $4.4$ $20.9$ $7.7$
Tax $13.5$ $10.5$ $26.5$ $15.9$ $24.1$
Sci+Com $25.2$ $12.8$ $12.5$ $27.2$ $25.3$
Tax+Com $22.1$ $8.0$ $19.5$ $25.1$ $30.0$
Mixture $26.8$ $24.7$ $26.4$ $30.0$ $30.5$
iNat21-2.7M Mixture $20.9$ $14.7$ $15.6$ $21.7$ $21.8$
ToL-10M Mixture $31.1$ $30.3$ $33.4$ $36.6$ $37.8$
Table 5: Zero-shot accuracy on species not seen during training (Rare Species task). Different rows and columns indicate different text types during training and test time, respectively. Blue indicates best accuracy and Orange indicates second-best. Using the taxonomic name over the scientific name always improves accuracy ($22.2\rightarrow$$26.5$ and $27.2\rightarrow$$30.0$). The final rows use the full iNat21 dataset and TreeOfLife-10M for reference.

### v3

@@ Abstract
Images of the natural world, collected by a variety of cameras, from drones to individual phones, are increasingly abundant sources of biological information. There is an explosion of computational methods and tools, particularly computer vision, for extracting biologically relevant information from images for science and conservation. Yet most of these are bespoke approaches designed for a specific task and are not easily adaptable or extendable to new questions, contexts, and datasets. A vision model for general organismal biology questions on images is of timely need. To approach this, we curate and release TreeOfLife-10M, the largest and most diverse ML-ready dataset of biology images. We then develop BioCLIP, a foundation model for the tree of life, leveraging the unique properties of biology captured by TreeOfLife-10M, namely the abundance and variety of images of plants, animals, and fungi, together with the availability of rich structured biological knowledge. We rigorously benchmark our approach on diverse fine-grained biology classification tasks and find that BioCLIP consistently and substantially outperforms existing baselines (by 16% to 17% absolute). Intrinsic evaluation reveals that BioCLIP has learned a hierarchical representation conforming to the tree of life, shedding light on its strong generalizability.111imageomics.github.io/bioclip has models, data and code.
@@ 1 Introduction
However, applying computer vision to answer a biological question is still a laborious task requiring substantial machine learning expertise and effort—biologists must manually label sufficient data for the specific taxa and task of interest, and find and train a suitable model for the task. Meanwhile, foundation models [12] such as CLIP [69] and GPT-3 [14] are extraordinarily valuable by enabling zero-shot or few-shot learning for a wide range of tasks. An analogous vision foundation model for biology should be useful for tasks spanning the entire tree of life [37, 53] instead of just the taxa it was trained on. Such a model would significantly lower the barrier to apply AI to biology.
@@
While the goals of generalization, fine-grained classification, and data efficiency are not new in computer vision, existing general-domain vision models [69, 95, 61] trained on hundreds of millions of images fall short when applied to evolutionary biology and ecology. Specifically, existing vision models produce general fine-grained representations, useful for comparing common organisms like dogs and wolves, but not for more fine-grained comparisons, e.g., Onoclea sensibilis and Onoclea hintonii (see Fig. 1).
@@
TreeOfLife-10M: a large-scale, diverse ML-ready biology image dataset. We curate and release the largest ML-ready dataset to-date of biology images with associated taxonomic labels, containing over $10$ million images covering $454$ thousand taxa in the tree of life. [^2] By ML-ready, we mean the data is standardized in a format suitable for training ML models and is readily available for downloading. In comparison, the current largest ML-ready biology image dataset, iNat21 [86], contains only $2.7$ million images covering $10$ thousand taxa. TreeOfLife-10M integrates existing high-quality datasets like iNat21 and Bioscan-1M [28]. More importantly, it includes newly curated images from the Encyclopedia of Life (eol.org), which supplies most of TreeOfLife-10M’s data diversity. Every image in TreeOfLife-10M is labeled with its taxonomic hierarchy to the finest level possible, as well as higher taxonomic ranks in the tree of life (see Figs. 1 and 3 for examples of taxonomic ranks and labels). TreeOfLife-10M enables training BioCLIP and future biology foundation models.
BioCLIP: a vision foundation model for the tree of life. With a large-scale labeled dataset like TreeOfLife-10M, a standard, intuitive training strategy (as adopted by other vision models like ResNet50 [33] and Swin Transformer [48]) is to use a supervised classification objective and learn to predict the taxonomic indices from an image. However, this fails to recognize and leverage the rich structure of taxonomic labels—taxa do not exist in isolation but are interconnected in a comprehensive taxonomy. Consequently, a model trained via plain supervised classification may not generalize well to taxa unseen in training, nor could it support zero-shot classification of unseen taxa.
@@
Comprehensive benchmarking. We comprehensively evaluate BioCLIP on $10$ fine-grained image classification datasets covering animals, plants, and fungi, including a newly curated Rare Species dataset unseen in training. BioCLIP achieves strong performance in both zero-shot and few-shot settings and substantially outperforms both CLIP [69] and OpenCLIP [42], leading to an average absolute improvement of 17% (zero-shot) and 16% (few-shot). Intrinsic analysis further reveals that BioCLIP has learned a more fine-grained hierarchical representation conforming to the tree of life, explaining its superior generalization.
@@ 2 TreeOfLife-10M
Recent work has shown that data quality and diversity are critical when training CLIP models [24, 57, 26]. We curate TreeOfLife-10M, the most diverse large-scale public ML-ready dataset for computer vision models in biology.
@@ 2.1 Images
The largest ML-ready biology image dataset is iNat21 [86] with 2.72.7M images of 1010K species. Despite this class breadth compared to popular general-domain datasets like ImageNet-1K [70], 1010K species is limited for biology. The International Union for Conservation of Nature (IUCN) reported over 22M total described species in 2022, with over 1010K bird species and over 1010K reptile species alone [44]. iNat21’s species diversity limits its potential for training a foundation model for the entire tree of life.
@@ 2.2 Metadata & Aggregation
The TreeOfLife-10M dataset integrates iNat21 (training split), our curated EOL dataset, and Bioscan-1M by aggregating the images and canonicalizing the labels. This is a highly non-trivial task because taxonomic hierarchies are notoriously noisy and rarely consistent between sources [36, 31, 52, 63, 4], likely contributing to the prior lack of image datasets large enough to train a foundation-scale vision model for the entire tree of life. We carefully unify and backfill taxonomic hierarchies from EOL, the Integrated Taxonomic Information System (ITIS) [43], and iNaturalist with special consideration for the existence of homonyms (genus-species labels shared among higher-order taxa). For more information on this process, the challenges, our solutions, and remaining issues, see Appendix C.
@@ 2.3 Release & Statistics
Our curated training and test datasets (TreeOfLife-10M and Rare Species, described in Sec. 4.2) are available on Hugging Face (with DOIs) under a public domain waiver, to the extent primary source licenses allow. This includes CSVs with image metadata and links to the primary sources, accompanied by a GitHub repository with the scripts to generate the datasets. [^4] We encourage future work to cite iNat21 [86], Bioscan-1M [28] and to appropriately attribute images from EOL based on their licenses if citing TreeOfLife-10M.
@@ 3.1 Why CLIP?
CLIP trains two uni-modal embedding models, a vision encoder and a text encoder, to (1) maximize feature similarity between positive (image, text) pairs and (2) minimize feature similarity between negative (image, text) pairs, where positive pairs are from the training data and negative pairs are all other possible (image, text) pairings in a batch. After training, CLIP’s encoder models embed individual instances of their respective modalities into a shared feature space. Next, we discuss formatting the text input to CLIP to incorporate the taxonomic structure.
@@ 4.1 Training and Evaluation Details
To train BioCLIP, we initialize from OpenAI’s CLIP weights [69] with a ViT-B/16 vision transformer [22] image encoder and a $77$-token causal autoregressive transformer text encoder. We continue pre-training on TreeOfLife-10M for 100 epochs with a cosine learning rate schedule [49]. We train on $8$ NVIDIA A100-80GB GPUs over $2$ nodes with a global batch size of $32,768$. We also train a baseline model on only the iNat21 dataset and multiple ablation models on 1M examples randomly sampled from TreeOfLife-10M (Secs. 4.3 and 4.4), following the same procedure for BioCLIP except with a smaller global batch size of $16,384$ on $4$ NVIDIA A100 GPUs on $1$ node. All hyperparameters and training details are in Appendix D and training and evaluation code is publicly available.
@@
We evaluate on 10 different classification tasks: the 8 biologically-relevant tasks from Meta-Album [84], Birds 525 [68] and our new Rare Species task (described in Sec. 4.2). Meta-Album is a dataset collection for meta-learning, encompassing various subjects. Specifically, we use the Plankton, Insects, Insects 2, PlantNet, Fungi, PlantVillage, Medicinal Leaf, and PlantDoc datasets. Our classification tasks cover all four multi-celled kingdoms in the tree of life (animals, plants, fungi, and protists) and have a diverse image distribution (photographs, microscope images, drawings, and museum specimens). Tab. 2 summarizes the datasets; they comprise a variety of label types from full taxonomic names to only scientific or common name.
@@
For zero-shot learning, we follow the same procedure as CLIP. For few-shot learning, we follow SimpleShot [90] and use a nearest-centroid classifier. For $k$-shot learning, we first randomly sample $k$ examples for each class and obtain the image embedding from the visual encoder of the pre-trained models. We then compute the average feature vector of the $k$ embeddings as the centroid for each class. All the examples left in the dataset are used for testing. After applying mean subtraction and L2-normalization to each centroid and test feature vector, we choose the class with the nearest centroid to the test vector as the prediction. We repeat each few-shot experiment $5$ times with different random seeds and report the mean accuracy in Tab. 4. Results with standard deviations are reported in Appendix E.
@@
We compare BioCLIP with the original OpenAI CLIP [69] and OpenCLIP [42] trained on LAION-400M [73]. Intuitively, common names of organisms are most pervasive in the training data of CLIP and OpenCLIP and these models work best with common names. This is also confirmed in our preliminary tests. Therefore, we use common names as class labels for CLIP and OpenCLIP by default unless unavailable for a dataset. BioCLIP can leverage taxonomic names, so we use taxonomic+common names by default. However, as noted in Tab. 2, the test datasets come in a variety of labels. Whenever the preferred label type is not available, we use labels that come with the dataset. We also compare to an ImageNet-21K [21] pre-trained model [78] and DINO [15] for few-shot classification.
@@
CLIP $49.9$ $3.2$ $9.1$ $9.8$ $58.5$ $10.2$ $5.4$ $15.9$ $26.1$ $31.8$ $21.9$
OpenCLIP $54.7$ $2.2$ $6.5$ $9.6$ $50.2$ $5.7$ $8.0$ $12.4$ $25.8$ $29.8$ $20.4$ 1.5-1.5
BioCLIP $72.1$ $6.1$ $34.8$ $20.4$ $91.4$ $40.7$ $24.4$ $38.6$ $28.4$ $38.0$ $39.4$ +17.5+17.5
– iNat21 Only $56.1$ $2.6$ $30.7$ $11.5$ $88.2$ $43.0$ $18.4$ $25.6$ $20.5$ $21.3$ $31.7$ +9.8+9.8
@@
CLIP $43.7$ $25.1$ $21.6$ $13.7$ $42.1$ $17.2$ $49.7$ $70.1$ $24.8$ $28.5$ $33.6$
OpenCLIP $53.7$ $32.3$ $23.2$ $14.3$ $45.1$ $18.4$ $53.6$ $71.2$ $26.8$ $29.2$ $36.7$ +3.1+3.1
Supervised-IN21K $60.2$ $22.9$ $14.7$ $14.4$ $46.7$ $16.9$ $62.3$ $58.6$ $27.7$ $28.0$ $35.2$ +1.6+1.6
DINO $40.5$ $37.0$ $23.5$ $16.4$ $30.7$ $20.0$ $60.0$ $79.2$ $23.7$ $31.0$ $36.2$ +2.6+2.6
BioCLIP $71.8$ $30.6$ $57.4$ $20.4$ $64.5$ $40.3$ $58.8$ $84.3$ $30.7$ $44.9$ $50.3$ +16.7+16.7
– iNat21 Only $74.8$ $29.6$ $53.9$ $19.7$ $67.4$ $35.5$ $55.2$ $75.1$ $27.8$ $36.9$ $47.5$ +13.9+13.9
@@
CLIP $73.5$ $41.2$ $39.9$ $24.6$ $65.2$ $27.9$ $71.8$ $89.7$ $35.2$ $46.0$ $51.5$
OpenCLIP $81.9$ $52.5$ $42.6$ $25.0$ $68.0$ $30.6$ $77.8$ $91.3$ $42.0$ $47.4$ $55.9$ +4.4+4.4
Supervised-IN21K $83.9$ $39.2$ $32.0$ $25.4$ $70.9$ $30.9$ $82.4$ $82.3$ $44.7$ $47.3$ $53.9$ +2.4+2.4
DINO $70.8$ $56.9$ $46.3$ $28.6$ $50.3$ $34.1$ $82.1$ $94.9$ $40.3$ $50.1$ $55.4$ +3.9+3.9
BioCLIP $90.0$ $49.3$ $77.8$ $33.6$ $85.6$ $62.3$ $80.9$ $95.9$ $47.5$ $65.7$ $68.8$ +17.3+17.3
– iNat21 Only $90.1$ $48.2$ $73.7$ $32.1$ $84.7$ $55.6$ $77.2$ $93.5$ $41.0$ $55.6$ $65.1$ +13.6+13.6
Table 4: Zero-, one- and five-shot classification top-1 accuracy for different models. Bold indicates best accuracy. All models use the same ViT-B/16 architecture. “iNat21 Only” follows the same procedure as BioCLIP but uses iNat21 instead of TreeOfLife-10M. $\Delta$ denotes the difference in mean accuracy with CLIP. Supervised-IN21K [78] and DINO [15] are vision-only models and cannot do zero-shot classification.
@@
ToL-1M Com $24.9$ $9.5$ $10.8$ $22.3$ $21.0$
Sci $11.0$ $22.3$ $4.5$ $21.5$ $8.0$
Tax $11.8$ $10.1$ $26.6$ $16.0$ $24.8$
Sci+Com $24.5$ $12.9$ $12.6$ $28.0$ $24.9$
Tax+Com $20.5$ $8.0$ $19.7$ $24.0$ $30.4$
Mixture $26.1$ $24.9$ $26.7$ $29.5$ $30.9$
iNat21-2.7M Mixture $20.4$ $14.7$ $15.6$ $20.9$ $21.3$
ToL-10M Mixture $31.6$ $30.1$ $34.1$ $37.0$ $38.0$
Table 5: Zero-shot accuracy on species not seen during training (Rare Species task). Different rows and columns indicate different text types during training and test time, respectively. Blue indicates best accuracy and Orange indicates second-best. Using the taxonomic name over the scientific name always improves accuracy ($22.3\rightarrow$$26.6$ and $28.0\rightarrow$$30.4$). The final rows use the full iNat21 dataset and TreeOfLife-10M for reference.